OpenStudy (khalilforthewin):

8x + 6y = 10 5x - 3y = 2.2 Which statement about the system of equations is true? A) The system has no solutions. B) The system has many solutions. C) The system has exactly 1 solution. D) The system has exactly 2 solutions.

OpenStudy (xmissalycatx):

Have you tried plugging any certain values into the equations to test it?

OpenStudy (khalilforthewin):

no i truely have no idea on how to find it @xMissAlyCatx

OpenStudy (xmissalycatx):

Okay, hold on.

OpenStudy (xmissalycatx):

Try to use this to guide you..

OpenStudy (khalilforthewin):

i got it, thanks

OpenStudy (etherealsarah):

In order to find y, we need to express x in terms of y , so we use the first equation to do that: 8x+6y=10 means that 8x=10-6y means x=(10-6y)/8 After that we head back towards the second equation and replace x: so 5x-3y=2.2 becomes 5(10-6y)/2 -3y=2.2 means 50-54y=17,6 means y=(17.6-50)/-54 means y=16.67 and then we return to the first equation where we expressed x in terms of y in order to find the value of x: x=(10-6x16.67)/8=-11.25 I am really tired so I could have committed a calculation error, you can check! And contact me if you do not understand anything! :)